# Einstein from Noise: Statistical Analysis

### v2

@@ H.1 The functional CLT for DFT
$\mathsf{Z}^{(d)}[k]\triangleq\frac{1}{\sqrt{d}}\sum_{\ell=0}^{d-1}z_{\ell}e^{-2\pi jk\ell/d},\quad 0\leq k<d.$
@@
$\mathsf{Z}^{(d)}=\big{(}\mathsf{Z}^{(d)}[0],\ldots,\mathsf{Z}^{(d)}[d-1],0,0,\ldots\big{)}\in\mathbb{C}^{\mathbb{N}}.$
@@
$\big{(}\mathsf{Z}^{(d)}[k_{1}],\ldots,\mathsf{Z}^{(d)}[k_{m}]\big{)}\xrightarrow[d\to\infty]{\mathcal{D}}\big{(}\mathsf{W}_{k_{1}},\ldots,\mathsf{W}_{k_{m}}\big{)},$ (H.1)
@@ H.2 Notations
$\mathsf{Z}^{(d)}[k]\triangleq\begin{cases}\displaystyle\frac{1}{\sqrt{d}}\sum_{\ell=0}^{d-1}z_{\ell}e^{-2\pi jk\ell/d},&0\leq k<d,\\ 0,&\text{otherwise}.\end{cases}$ (H.2)
@@
$\hat{\mathsf{R}}_{\mathsf{Z}}^{(d)}\triangleq\operatorname*{arg\,max}_{0\leq r<d}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Z}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Z}}[k]-\phi_{\mathsf{X}}[k]\right),$ (H.3)
@@
$\phi_{e,\mathsf{Z}}^{(d)}[k]\triangleq\frac{2\pi k\hat{\mathsf{R}}_{\mathsf{Z}}^{(d)}}{d}+\phi_{\mathsf{Z}}[k]-\phi_{\mathsf{X}}[k].$ (H.4)
@@
$\mathsf{N}^{(d)}[k]\triangleq\begin{cases}\displaystyle\frac{1}{\sqrt{d}}\sum_{\ell=0}^{d-1}n_{\ell}e^{-2\pi jk\ell/d},&0\leq k<d,\\ 0,&\text{otherwise}.\end{cases}$ (H.5)
@@
$\hat{\mathsf{R}}_{\mathsf{N}}^{(d)}\triangleq\operatorname*{arg\,max}_{0\leq r<d}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right),$ (H.6)
@@
$\phi_{e,\mathsf{N}}^{(d)}[k]\triangleq\frac{2\pi k\hat{\mathsf{R}}_{\mathsf{N}}^{(d)}}{d}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k].$ (H.7)
H.3 Convergence of the real and imaginary parts of the EfN estimator
We now present an auxiliary result that relates the real and imaginary parts of the EfN estimator under both Gaussian i.i.d. and general i.i.d. noise models. This result is a consequence of the functional central limit theorem for the DFT (Theorem H.1).
Proposition H.2.
Let $z_{0},z_{1},\ldots$ be a sequence of i.i.d. real-valued random variables with zero mean, finite variance, $\mathbb{E}[z_{0}^{2}]=\sigma^{2}<\infty$, and finite fourth moment. Let $n_{0},n_{1},\ldots$ be an i.i.d. sequence of Gaussian random variables with $n_{\ell}\sim\mathcal{N}(0,\sigma^{2})$. Let $\mathsf{Z}^{(d)}$ and $\mathsf{N}^{(d)}$ denote the DFTs of the sequences $\{z_{i}\}_{i=0}^{d-1}$ and $\{n_{i}\}_{i=0}^{d-1}$, respectively. Let $\phi_{e,\mathsf{Z}}^{(d)}[k]$ and $\phi_{e,\mathsf{N}}^{(d)}[k]$ denote the aligned phase differences defined in (H.4) and (H.7), respectively. Then, for each fixed frequency index $k\in\mathbb{N}$,
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0,$ (H.8)
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0,$ (H.9)
@@ H.3 Convergence of the real and imaginary parts of the EfN estimator
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0,$ (H.10)
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0.$ (H.11)
Proof of Proposition H.2.
We begin by applying Theorem H.1 to the sequence $\{z_{i}\}$, which satisfies its assumptions. This in turn implies that,
$\mathsf{Z}^{(d)}\xrightarrow{\mathcal{D}}\mathsf{W},$ (H.12)
@@
$\mathsf{N}^{(d)}\overset{\mathcal{D}}{=}(\mathsf{W}[0],\ldots,\mathsf{W}[d-1]),$ (H.13)
@@
$\mathbb{E}[|\mathsf{Z}^{(d)}[k]|^{2}]<\infty,$ (H.14)
@@
$\left\{|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},\quad\left\{|\mathsf{Z}^{(d)}[k]|\cos(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},$ (H.15)
@@
$\lim_{d\to\infty}\biggl{(}\mathbb{E}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathbb{E}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right]\biggr{)}=0.$ (H.16)
@@
$\mathbb{E}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]=\mathbb{E}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right],\quad\forall d\in\mathbb{N}^{+}.$ (H.17)
Combining (H.16) and (H.17) yields
$\lim_{d\to\infty}\biggl{(}\mathbb{E}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathbb{E}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]\biggr{)}=0,$ (H.18)
which establishes (H.8). An identical argument with sine replaced by cosine proves (H.9).
@@
$\mathbb{E}[|\mathsf{Z}^{(d)}[k]|^{4}]<\infty,$ (H.19)
@@
$\left\{|\mathsf{Z}^{(d)}[k]|^{2}\sin^{2}(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},\quad\left\{|\mathsf{Z}^{(d)}[k]|^{2}\cos^{2}(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},$ (H.20)
@@
$\lim_{d\to\infty}\biggl{(}\mathsf{Var}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathsf{Var}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right]\biggr{)}=0.$ (H.21)
@@
$\mathsf{Var}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]=\mathsf{Var}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right],\quad\forall d\in\mathbb{N}^{+}.$ (H.22)
Combining (H.21) and (H.22) yields,
$\lim_{d\to\infty}\biggl{(}\mathsf{Var}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathsf{Var}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]\biggr{)}=0,$ (H.23)
establishing (H.10). Repeating the same steps with cosine in place of sine proves (H.11).
H.4 Proof of Theorem 5.2
@@
$\mathsf{\hat{X}}[k]e^{-j\phi_{\mathsf{X}}[k]}=$ $\frac{1}{M}{\sum_{i=0}^{M-1}\left|\mathsf{Z}_{i}[k]\right|e^{j\phi_{e,\mathsf{Z}_{i}}[k]}}$ (H.24)
${\mathsf{a.s.}}\mathbb{E}\left[\left|\mathsf{Z}_{1}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{Z}_{1}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}[k]\right)\right],$ (H.25)
@@
$\phi_{e,\mathsf{Z}_{i}}^{(d)}[k]\triangleq\frac{2\pi k\hat{\mathsf{R}}_{\mathsf{Z}_{i}}^{(d)}}{d}+\phi_{\mathsf{Z}_{i}}[k]-\phi_{\mathsf{X}}[k].$ (H.26)
@@
$\hat{\mathsf{R}}_{\mathsf{Z}_{i}}^{(d)}\triangleq\operatorname*{arg\,max}_{0\leq r<d}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Z}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Z}_{i}}[k]-\phi_{\mathsf{X}}[k]\right).$ (H.27)
Next, we invoke Proposition H.2, whose assumptions are satisfied in this setting. Let $\mathsf{N}_{1}^{(d)}$ denotes a noise vector with i.i.d. Gaussian entries that match the first and second moments of the entries of $\mathsf{Z}_{1}^{(d)}$ (as defined in Proposition H.2). We note that the results of Theorems 4.1 and 4.3 apply to the case of i.i.d. Gaussian entries $\mathsf{N}_{1}^{(d)}$. Then, by Proposition H.2, for each fixed frequency index $k\in\mathbb{N}$, the following convergence results hold,
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0,$ (H.28)
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0.$ (H.29)
@@
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0,$ (H.30)
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0.$ (H.31)
By Theorems 4.1 and 4.3, the convergence behavior of the estimator is governed by the variances in (H.30) and (H.31). Therefore, the asymptotic behavior of the estimator for general i.i.d. noise $\{z_{i}\}$ matches that of the Gaussian i.i.d. case $\{n_{i}\}$. In particular for (5.3), we have,
$\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]$ $=\arctan\left(\frac{\sum_{i=0}^{M-1}\left|\mathsf{Z}_{i}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{i}}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Z}_{i}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{i}}[k]\right)}\right)$ (H.32)
[…]
where (H.33) follows from the SLLN as $M\to\infty$. Applying (H.28) and (H.29) into (H.33), yields,
$\lim_{d\to\infty}\lim_{M\to\infty}\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=\lim_{d\to\infty}\arctan\left(\frac{\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}_{1}}[k]\right)\right]}{\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}_{1}}[k]\right)\right]}\right).$ (H.34)
By (C.6), the r.h.s. of (H.34) vanishes for every $d$, and therefore,
$\lim_{d\to\infty}\lim_{M\to\infty}\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=0,$ (H.35)

### v3

@@ F.1 The functional CLT for DFT
$\mathsf{Z}^{(d)}[k]\triangleq\frac{1}{\sqrt{d}}\sum_{\ell=0}^{d-1}z_{\ell}e^{-2\pi jk\ell/d},\quad 0\leq k<d.$ (F.1)
@@
$\mathsf{Z}^{(d)}=\big(\mathsf{Z}^{(d)}[0],\ldots,\mathsf{Z}^{(d)}[d-1],0,0,\ldots\big)\in\mathbb{C}^{\mathbb{N}}.$ (F.2)
@@
$\big(\mathsf{Z}^{(d)}[k_{1}],\ldots,\mathsf{Z}^{(d)}[k_{m}]\big)\xrightarrow[d\to\infty]{\mathcal{D}}\big(\mathsf{W}_{k_{1}},\ldots,\mathsf{W}_{k_{m}}\big),$ (F.3)
@@ F.2 Notations
$\mathsf{Z}^{(d)}[k]\triangleq\begin{cases}\displaystyle\frac{1}{\sqrt{d}}\sum_{\ell=0}^{d-1}z_{\ell}e^{-2\pi jk\ell/d},&0\leq k<d,\\ 0,&\text{otherwise}.\end{cases}$ (F.4)
@@
$\hat{\mathsf{R}}_{\mathsf{Z}}^{(d)}\triangleq\operatorname*{arg\,max}_{0\leq r<d}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Z}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Z}}[k]-\phi_{\mathsf{X}}[k]\right),$ (F.5)
@@
$\phi_{e,\mathsf{Z}}^{(d)}[k]\triangleq\frac{2\pi k\hat{\mathsf{R}}_{\mathsf{Z}}^{(d)}}{d}+\phi_{\mathsf{Z}}[k]-\phi_{\mathsf{X}}[k].$ (F.6)
@@
$\mathsf{N}^{(d)}[k]\triangleq\begin{cases}\displaystyle\frac{1}{\sqrt{d}}\sum_{\ell=0}^{d-1}n_{\ell}e^{-2\pi jk\ell/d},&0\leq k<d,\\ 0,&\text{otherwise}.\end{cases}$ (F.7)
@@
$\hat{\mathsf{R}}_{\mathsf{N}}^{(d)}\triangleq\operatorname*{arg\,max}_{0\leq r<d}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right),$ (F.8)
@@
$\phi_{e,\mathsf{N}}^{(d)}[k]\triangleq\frac{2\pi k\hat{\mathsf{R}}_{\mathsf{N}}^{(d)}}{d}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k].$ (F.9)
F.3 Convergence of the real and imaginary parts of the EfN estimator
We now present an auxiliary result that relates the real and imaginary parts of the EfN estimator under both Gaussian i.i.d. and general i.i.d. noise models. This result is a consequence of the functional central limit theorem for the DFT (Theorem F.1).
Proposition F.2.
Let $z_{0},z_{1},\ldots$ be a sequence of i.i.d. real-valued random variables with zero mean, finite variance, $\mathbb{E}[z_{0}^{2}]=\sigma^{2}<\infty$, and finite fourth moment. Let $n_{0},n_{1},\ldots$ be an i.i.d. sequence of Gaussian random variables with $n_{\ell}\sim\mathcal{N}(0,\sigma^{2})$. Let $\mathsf{Z}^{(d)}$ and $\mathsf{N}^{(d)}$ denote the DFTs of the sequences $\{z_{i}\}_{i=0}^{d-1}$ and $\{n_{i}\}_{i=0}^{d-1}$, respectively. Let $\phi_{e,\mathsf{Z}}^{(d)}[k]$ and $\phi_{e,\mathsf{N}}^{(d)}[k]$ denote the aligned phase differences defined in (F.6) and (F.9), respectively. Then, for each fixed frequency index $k\in\mathbb{N}$,
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0,$ (F.10)
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0,$ (F.11)
@@ F.3 Convergence of the real and imaginary parts of the EfN estimator
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0,$ (F.12)
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}}^{(d)}[k]\right)\right]\right)=0.$ (F.13)
Proof of Proposition F.2.
We begin by applying Theorem F.1 to the sequence $\{z_{i}\}$, which satisfies its assumptions. This in turn implies that,
$\mathsf{Z}^{(d)}\xrightarrow{\mathcal{D}}\mathsf{W},$ (F.14)
@@
$\mathsf{N}^{(d)}\overset{\mathcal{D}}{=}(\mathsf{W}[0],\ldots,\mathsf{W}[d-1]),$ (F.15)
@@
$\mathbb{E}[|\mathsf{Z}^{(d)}[k]|^{2}]<\infty,$ (F.16)
@@
$\left\{|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},\quad\left\{|\mathsf{Z}^{(d)}[k]|\cos(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},$ (F.17)
@@
$\lim_{d\to\infty}\biggl(\mathbb{E}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathbb{E}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right]\biggr)=0.$ (F.18)
@@
$\mathbb{E}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]=\mathbb{E}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right],\quad\forall d\in\mathbb{N}^{+}.$ (F.19)
Combining (F.18) and (F.19) yields
$\lim_{d\to\infty}\biggl(\mathbb{E}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathbb{E}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]\biggr)=0,$ (F.20)
which establishes (F.10). An identical argument with sine replaced by cosine proves (F.11).
@@
$\mathbb{E}[|\mathsf{Z}^{(d)}[k]|^{4}]<\infty,$ (F.21)
@@
$\left\{|\mathsf{Z}^{(d)}[k]|^{2}\sin^{2}(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},\quad\left\{|\mathsf{Z}^{(d)}[k]|^{2}\cos^{2}(\phi_{e,\mathsf{Z}}^{(d)}[k])\right\}_{d\in\mathbb{N}^{+}},$ (F.22)
@@
$\lim_{d\to\infty}\biggl(\mathsf{Var}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathsf{Var}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right]\biggr)=0.$ (F.23)
@@
$\mathsf{Var}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]=\mathsf{Var}\left[|\mathsf{W}[k]|\sin(\phi_{e,\mathsf{W}}[k])\right],\quad\forall d\in\mathbb{N}^{+}.$ (F.24)
Combining (F.23) and (F.24) yields,
$\lim_{d\to\infty}\biggl(\mathsf{Var}\left[|\mathsf{Z}^{(d)}[k]|\sin(\phi_{e,\mathsf{Z}}^{(d)}[k])\right]-\mathsf{Var}\left[|\mathsf{N}^{(d)}[k]|\sin(\phi_{e,\mathsf{N}}^{(d)}[k])\right]\biggr)=0,$ (F.25)
establishing (F.12). Repeating the same steps with cosine in place of sine proves (F.13).
F.4 Proof of Theorem 5.2
@@
$\mathsf{\hat{X}}[k]e^{-j\phi_{\mathsf{X}}[k]}=$ $\frac{1}{M}{\sum_{i=0}^{M-1}\left|\mathsf{Z}_{i}[k]\right|e^{j\phi_{e,\mathsf{Z}_{i}}[k]}}$ (F.26)
$\xrightarrow[]{\mathsf{a.s.}}\mathbb{E}\left[\left|\mathsf{Z}_{1}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}[k]\right)\right]+j\mathbb{E}\left[\left|\mathsf{Z}_{1}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}[k]\right)\right],$ (F.27)
@@
$\phi_{e,\mathsf{Z}_{i}}^{(d)}[k]\triangleq\frac{2\pi k\hat{\mathsf{R}}_{\mathsf{Z}_{i}}^{(d)}}{d}+\phi_{\mathsf{Z}_{i}}[k]-\phi_{\mathsf{X}}[k].$ (F.28)
@@
$\hat{\mathsf{R}}_{\mathsf{Z}_{i}}^{(d)}\triangleq\operatorname*{arg\,max}_{0\leq r<d}\sum_{k=0}^{d-1}\left|\mathsf{X}[k]\right|\left|\mathsf{Z}_{i}[k]\right|\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Z}_{i}}[k]-\phi_{\mathsf{X}}[k]\right).$ (F.29)
Next, we invoke Proposition F.2, whose assumptions are satisfied in this setting. Let $\mathsf{N}_{1}^{(d)}$ denotes a noise vector with i.i.d. Gaussian entries that match the first and second moments of the entries of $\mathsf{Z}_{1}^{(d)}$ (as defined in Proposition F.2). We note that the results of Theorems 4.1 and 4.3 apply to the case of i.i.d. Gaussian entries $\mathsf{N}_{1}^{(d)}$. Then, by Proposition F.2, for each fixed frequency index $k\in\mathbb{N}$, the following convergence results hold,
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0,$ (F.30)
$\lim_{d\to\infty}\left(\mathbb{E}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0.$ (F.31)
@@
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0,$ (F.32)
$\lim_{d\to\infty}\left(\mathsf{Var}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}^{(d)}[k]\right)\right]-\mathsf{Var}\left[\left|\mathsf{N}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}_{1}}^{(d)}[k]\right)\right]\right)=0.$ (F.33)
By Theorems 4.1 and 4.3, the convergence behavior of the estimator is governed by the variances in (F.32) and (F.33). Therefore, the asymptotic behavior of the estimator for general i.i.d. noise $\{z_{i}\}$ matches that of the Gaussian i.i.d. case $\{n_{i}\}$. In particular for (5.3), we have,
$\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]$ $=\arctan\left(\frac{\sum_{i=0}^{M-1}\left|\mathsf{Z}_{i}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{i}}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Z}_{i}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{i}}[k]\right)}\right)$ (F.34)
$\xrightarrow[]{\text{a.s.}}\arctan\left(\frac{\mathbb{E}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{Z}_{1}}[k]\right)\right]}{\mathbb{E}\left[\left|\mathsf{Z}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{Z}_{1}}[k]\right)\right]}\right),$ (F.35)
where (F.35) follows from the SLLN as $M\to\infty$. Applying (F.30) and (F.31) into (F.35), yields,
$\lim_{d\to\infty}\lim_{M\to\infty}\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=\lim_{d\to\infty}\arctan\left(\frac{\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\sin\left(\phi_{e,\mathsf{N}_{1}}[k]\right)\right]}{\mathbb{E}\left[\left|\mathsf{N}_{1}^{(d)}[k]\right|\cos\left(\phi_{e,\mathsf{N}_{1}}[k]\right)\right]}\right).$ (F.36)
By (B.62), the r.h.s. of (F.36) vanishes for every $d$, and therefore,
$\lim_{d\to\infty}\lim_{M\to\infty}\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=0,$ (F.37)